OpenStudy (anonymous):

how would i get this into y=mx+b form? 5x-2y=-16

OpenStudy (anonymous):

let y be the subject, find y and then you'd get to compare your new equation with y=mx+c.

OpenStudy (anonymous):

5x-2y=-16 5x+16=2y 5/2x+8=y

OpenStudy (anonymous):

saif u are wrong

OpenStudy (saifoo.khan):

Oh right, my bad. sorry!

OpenStudy (anonymous):

yeah i knew he was wrong but I figured out the answer after but thanks guys!

OpenStudy (anonymous):

\[y = \frac{5}{2}x + 8\]

OpenStudy (anonymous):

\[\huge{\begin{array}l\color{#FF0000}{\text{l}}\color{#FF7F00}{\text{I}}\color{#FFFF00}{\text{I}}\color{#00FF00}{\text{l}}\color{#0000FF}{\text{I}}\color{#6600FF}{\text{I}}\color{#8B00FF}{\text{I}}\color{#FF0000}{\text{l}}\color{#FF7F00}{\text{I}}\color{#FFFF00}{\text{l}}\color{#00FF00}{\text{l}}\color{#0000FF}{\text{I}}\color{#6600FF}{\text{I}}\color{#8B00FF}{\text{I}}\color{#FF0000}{\text{I}}\color{#FF7F00}{\text{l}}\color{#FFFF00}{\text{l}}\color{#00FF00}{\text{l}}\color{#0000FF}{\text{I}}\color{#6600FF}{\text{l}}\color{#8B00FF}{\text{I}}\color{#FF0000}{\text{l}}\color{#FF7F00}{\text{I}}\end{array}}\]